A car travels 1km it starts to accelerate at 2.5m/s it attains a velocity of 12.5m/s the car continues at this velocity for some time and decelerates at 3m/s^2 until it stops. What is the total time for the whole distance?


1
Expert's answer
2020-08-20T09:48:47-0400

Let the first acceleration be "a_1 = 2.5 m\/s^2" and the second one "a_2 = 3m\/s^2". The Initial speed is "v_0 = 0", the next one is "v_1 = 12.5m\/s" and the final is "v_2 = 0". The total distance is "d = 1km = 1000m".

Let's find the time "t_1" of the first acceleration from "v_0" to "v_1"

By definition, the acceleration is:


"a_1 = \dfrac{v_1 - v_0}{t_1}\\
t_1 = \dfrac{v_1 - v_0}{a_1} = \dfrac{12.5-0}{2.5} = 5s"

During this time the car covered the distance:


"d_1 = \dfrac{a_1t_1^2}{2} = \dfrac{2.5\cdot 5^2}{2} = 31.25 m"

Let's find the time "t_3" of the last acceleration. Again, by definition:


"a_2 = \dfrac{v_2-v_1}{t_3}\\
t_1 = \dfrac{|v_2 - v_1|}{a_2} = \dfrac{|0-12.5|}{3} = \dfrac{25}{6}s"

During this time the car covered the distance:


"d_3 = \dfrac{a_2t_3^2}{2} = \dfrac{3\cdot 25^2}{2\cdot 6^2} = 156.25 m"

The distance in the middle, covered with a constant speed "v_1" is then:


"d_2 = d-d_1 - d_3 = 1000m - 31.25m - 156.25m = 812.5m"

Thus, the time during which it was covered is:


"t_2 = d_2\/v_1 = 65 s"

Thus, the total time is:


"t = t_1 + t_2 + t_3 = 5s + \dfrac{25}{6}s + 65s \approx 74.17s"

Answer. 74.17s.